# Symmetries and singular behaviors with Bohmian trajectories

### v1

@@ Abstract
Quantum mechanics is able to produce challenging behaviors even in the simplest situations. This is possible because of the important dynamical role that phase plays in the evolution of quantum systems, very similar, nonetheless, to effects observable in analogous optical systems. The present work focuses on how Bohmian mechanics proves to be a rather convenient theoretical framework to analyze phase-based phenomena, since the phase constitutes the central element in this hydrodynamical formulation of quantum mechanics. More specifically, it allows us to understand how spatial phase variations give rise to velocity fields that eventually rule the dynamical behavior of quantum systems, and that, when integrated in time locally (i.e., taking into account specific positions), they provide us with a neat local (point by point) description of the system evolution in the configuration space. Indeed, it will also be seen that this idea transcends the quantum realm and can be profitably used to describe the behavior of optical analogs with rather singular behaviors. With this purpose, two interesting phenomena that take place in free space are considered, namely, the self-acceleration and shape-invariance of Airy beams, and spontaneous self-focusing.
@@ 1 Introduction
Since its formulation by David Bohm in 1952 [1, 2], Bohmian mechanics has been widely understood as a causal interpretation or model for quantum phenomena, alternative to the current quantum theory. With its seminal work, Bohm opened up a new way to conceive quantum physics, aligned with other alternative approaches aimed at favoring a quantum theory without observers [3, 4]. Yet, Bohmian mechanics can also be simply considered, from a strictly formal point of view, a direct consequence of reformulating standard quantum mechanics in hydrodynamic terms, as it was done by Erwin Madelung in 1926 [5], which also leads to interesting implications and applications [6, 7, 8, 9, 10]. In this case, the concept of probability current or quantum flux [11, 12] acquires a major role, since it constitutes the direct link connecting the usual evolution of the probability density in the configuration space, described by the Schrödinger equation, with the so-called Bohmian momentum or guidance condition. Thus, it is not that Bohm’s momentum is any external concepts to the usual quantum theory, but it is contained within it [13]. Indeed, physically, it is a measure of the local spatial variations undergone by the phase, while the so-called Bohmian trajectories constitute, in turn, a local measure of the flux along precise streamlines, which have nothing to do with real trajectories followed by a quantum system [14, 15, 16].
@@
Therefore, we find that the so-called Bohmian trajectories can be understood either as hidden variables, providing us with an alternative perspective of the quantum realm [17, 18, 19, 20, 6, 21, 22, 23], or as a methodological tool, worth applying to a wide variety of physical problems to get new insights [7, 8, 9, 24, 10], which can have either a quantum mechanical nature, or necessarily not. Here, we are going to focus on the second point of view, trying to settle in simple terms what a Bohmian trajectory is [15, 16] and what it can be good for [13], leaving aside expendable ontological connotations. This will also serve us to show how this quantum trajectory-based methodology can be extrapolated to other non-quantum contexts, which keep a certain resemblance at a formal level.
@@
With that in mind, two challenging wave phenomena in free space, directly associated with the phase, are going to be analyzed by means of Bohmian mechanics: (1) solutions that contain an infinite amount of energy, with particular emphasis on the phenomenology of Airy wave packets due to their self-acceleration and shape-invariance properties [25], and (2) self-focusing solutions [26], i.e., solutions that converge to a singularity with time without the action of any external agent or nonlinear contribution [27, 28, 29, 30]. In both cases, the discussion will show how, despite of the peculiar nature of these solutions, they admit well-posed trajectory-based descriptions, which help us to understand in an intuitive manner the main features (the challenging ones) of their respective dynamical evolutions and, from them, the major role played in those dynamics by the local variations of the quantum phase. An interesting point here is thus the one concerning the symmetries displayed by the underlying velocity fields (or, equivalently, the associated phase fields), and hence the physical implications arising when such symmetries are broken.
@@
The work is organized as follows. In next section, Sec. 2, a brief overview of the main theoretical aspects related to Bohmian mechanics is summarized, as well as its extension to standard paraxial wave optics, which relies on a Schrödinger-type equation, namely, the paraxial Helmholtz equation. Section 3 presents a discussion of the Airy beam propagation in terms of Bohmian trajectories, while spontaneous self-focusing is considered in Sec. 4. Finally, a series of concluding remarks are highlighted in Sec. 5.
@@ 2.1 Velocities and trajectories in standard quantum mechanics
From an essentially pragmatic point of view, i.e., leaving aside ontological issues, Bohmian trajectories from a pragmatic point of view, that is, simply following standard quantum mechanics. Thus, consider a spinless, nonzero mass quantum system described in the configuration space, at a time $t=0$, by the wave function $\psi({\bf r},0)$. If the system is in free space, the time evolution of the system wave function at any other later time is ruled by the simple time-dependent Schrödinger equation
@@
${\bf j}({\bf r},t)=\frac{\hbar}{2mi}\left[\psi^{*}({\bf r},t)\nabla\psi({\bf r},t)-\psi({\bf r},t)\nabla\psi^{*}({\bf r},t)\right]$ (3)
@@
Note here, in the last equality, that $\hat{\bf p}=-i\hbar\nabla$ is the usual momentum operator, and hence ${\bf p}({\bf r},t)=m{\bf v}({\bf r},t)$ can be understood as a measure of the local value of the momentum, complementary of the corresponding expectation value, which provides us with a sort of average of ${\bf p}$ over the whole configuration space.
So far, everything is included within standard quantum mechanics, and therefore there is no need to postulate anything. The conventional Bohmian approach merges standard theory with the concept of hidden variable, which is defined as a specific position in configuration space [1, 2, 6, 31]. The evolution in time of this position generates a formally well-defined trajectory, which is what we typically call a Bohmian trajectory, and that does not violate any fundamental law of quantum mechanics, such as the uncertainty relations or the complementarity principle. To determine such trajectories, Bohm considered [1, 6] the same stating point as here, with the additional nonlinear transformation relation $\{\psi,\psi^{*}\}\to\{\rho,S\}$, from two complex field variables to two real field variables, such that
@@
and the same for $\psi^{*}$. After making the corresponding substitution into Eq. (1), and its conjugate, and separating the real and imaginary parts of the resulting equation, he reached the equations of motion for both $\rho({\bf r},t)$ and $S({\bf r},t)$, namely,
$\frac{\partial\rho({\bf r},t)}{\partial t}$ $=$ $-\nabla\left[\rho({\bf r},t)\frac{\nabla S({\bf r},t)}{m}\right],$ (6)
@@
Equation (6) is the continuity equation (2), although the velocity field is recast in terms of the new phase variable, ${\bf v}=\nabla S/m$, while Eq. (7) is the so-called quantum Hamilton-Jacobi equation, which generalizes the classical version [6, 32] by considering the extra term in it $Q({\bf r},t)$, known as Bohm’s potential or quantum potential,
@@
$\dot{\bf r}({\bf r},t)=\frac{{\bf p}({\bf r},t)}{m}=\frac{\nabla S({\bf r},t)}{m}$ (9)
where ${\bf p}$ is the so-called Bohm’s momentum, and this equation of motion is known as the guidance condition or equation. Leaving aside the discussion on the existence of hidden variables, the fact is that this equation is well-posed mathematically and, therefore, there is not reason for it to be rejected, other than ill-founded prejudices based on the interpretation that some might do of its outcomes, the trajectories. From a formal point of view, though, this equation simply tells us that what governs how the probability density distributes in configuration space along time is an underlying velocity field associated with local spatial variations of the system’s phase. Maybe this is a rather challenging consequence in quantum mechanics. However, note that in optics the engineering of structured light beams [33], which precisely follow similar equations (see below), is actually possible by making light to pass through devices that produce such spatial phase variations (i.e., spatial light modulators). Even more, with the possibility to experimentally obtain weak measurements [34, 35], it has been possible to measure this transverse field and, from it, to infer the path followed by the corresponding trajectories [36, 37, 38] (which has nothing to do with the hypothetical pathways that could be expected to be followed by individual photons).
From a conventional point of view, we thus have a combination of standard tools in a hydrodynamic language (à la Madelung [5]) and a postulated momentum (à la Bohm [1]), which does not disprove the above pragmatic consideration of Bohmian mechanics as a valid methodology or quantum representation, beyond its commonly accepted conception of being a hidden-variable model. This leads us to a third aspect, which emphasizes the usefulness of this approach as a convenient computational tool. Specifically, when one deals with the issue of propagating the wave function following the Schrödinger equation, there are different strategies, the so-called spectral methods being a widely used methodology [39, 40, 41]. Basically, the idea in these methods is to recast the wave function as a linear superposition of eigenfunctions of the Laplacian, so that the nonlocal action of the kinetic term can be evaluated in a more accurate manner than with discretizing methods (where the accuracy of the method is associated with the number of neighbors included in the approximation of the discrete Laplacian). From a Bohmian point of view, this is a rather convenient strategy, for one can obtain analytical expressions for the equation of motion in terms of the spectral or pseudo-spectral components of the wave function [9]. Indeed, if the method is based on the Fourier transform, these components can directly be the components of the discrete Fourier transform, which simplifies remarkably the numerical calculation of the trajectories and, at the same time, also increases their numerical accuracy (something very important in problems characterized by the presence and sudden emergence of nodes).
@@
where $c_{j}=|c_{j}|e^{\phi_{j}}$, ${\bf k}_{j}={\bf p}_{j}/\hbar$, and $E_{j}={\bf p}_{j}^{2}/2m$. Note that the ${\bf k}_{j}$ might correspond, for instance, to the discrete set of momenta that arise as soon as the configuration space is discretized in a mess of equidistant knots, when we apply the fast Fourier method [42]. If Eq. (10) is substituted into Eqs. (4) or, equivalently, (9), we obtain the equation of motion
@@
2.2 Velocities and trajectories in standard quantum mechanics
@@
Let us now assume that the energy propagates mostly along a given direction (say, $z$), while the effects of interest lie on perpendicular planes. This means that we can assume paraxial conditions, i.e., all important effects take place close the $z$-axis as the field mainly moves forward parallel to it. Accordingly, we can write the time-independent field amplitude as $\Psi({\bf r}_{\perp},z)\approx\psi({\bf r}_{\perp},z)e^{ikz}$, i.e., as a plane wave accounting for the propagation along the $z$-axis, with a modulating amplitude along the transverse direction. With this assumption, and after neglecting slow second order derivatives, Eq. (15) becomes the paraxial Helmholtz equation,
@@
and $\nabla^{2}_{\perp}$ is replaced by $\nabla^{2}$, then we immediately recover Schrödinger’s equation for propagation in free space. Accordingly, considering the same transformation, from Eq. (9), we find that an analogous trajectory equation can also be obtained in optics, since
@@
Note that the above relation basically says that the momentum and the wave vector are linearly related one another, since, making both sides to be equal, we obtain ${\bf p}/\hbar={\bf k}$, which is precisely de Broglie’s quantization condition, which dynamically associates the particle aspect (${\bf p}$) with a wave aspect (${\bf k}$ or, equivalently, $1/\lambda$) — also valid for photons if Einstein’s equations for the energy of a zero-mass particle and for the photon energy are made equal.
@@
where ${\bf v}_{\perp}$ is not a velocity in strict sense, but a field that behaves similarly in an “effective” manner. This effective field is analogous to the transverse momentum introduced in [45] to describe the properties of an atomic Mach-Zehnder interferometer. The integration of this quantity in $z$ provides us with trajectories that describe the spatial distribution of the electromagnetic energy or, equivalently, how the latter redistributes along the transverse direction as $z$ increases. It is worth noting that the outcomes experimentally reported in [37] correspond precisely to the transverse field (19), while the inferred trajectories are in compliance with what one would obtain by integrating this effective equation of motion. On the other, it is also worth emphasizing that Eq. (19) for paraxial conditions is an approximation to the general expression that arises from the Poynting vector [46].
@@
where ${\bf u}$ refers to the spatial coordinates (transverse, in the case of the paraxial optical system), $\xi$ is the propagation variable, and $\nabla^{2}_{\bf u}$ and $\nabla_{\bf u}$ are, respectively, the Laplacian and the gradient computed with respect to the spatial coordinates ${\bf u}$, and ${\bf v}_{\bf u}$ denotes the effective velocity field in the space defined by the coordinates ${\bf u}$.

### v2

@@ Abstract
Quantum mechanics is able to predict challenging behaviors even in the simplest physical scenarios. These behaviors are possible because of the important dynamical role that phase plays in the evolution of quantum systems, and are very similar, on the other hand, to effects observable in analogous optical systems. This work focuses on how Bohmian mechanics proves to be a rather convenient theoretical framework to analyze phase-based phenomena, since the phase constitutes the central element in this hydrodynamical formulation of quantum mechanics. More specifically, it allows us to understand how spatial phase variations give rise to velocity fields that eventually rule the dynamical behavior of quantum systems, and that, when integrated in time locally (i.e., taking into account specific positions), they provide us with a neat local (point by point) description of the system evolution in the configuration space. Indeed, it will also be seen that this idea transcends the quantum realm and can be profitably used to describe the behavior of optical analogs with rather singular behaviors. With this purpose, two interesting phenomena that take place in free space are considered, namely, the self-acceleration and shape-invariance of Airy beams, and spontaneous self-focusing.
@@ 1 Introduction
Since its formulation by David Bohm in 1952 [1, 2], Bohmian mechanics has been widely understood as a causal interpretation or model for quantum phenomena, alternative to the current quantum theory. With its seminal work, Bohm opened up a new way to conceive quantum physics, aligned with other alternative approaches aimed at favoring a quantum theory without observers [3, 4]. Yet, Bohmian mechanics can also be simply considered, from a strictly formal point of view, a direct consequence of reformulating standard quantum mechanics in hydrodynamic terms, as it was done by Erwin Madelung in 1926 [5], which also leads to interesting implications and applications [6, 7, 8, 9, 10]. In this case, the concept of probability current or quantum flux [11, 12] acquires a major role, since it constitutes a direct link that connects the usual evolution of the probability density in the configuration space, described by the Schrödinger equation, with the so-called Bohmian momentum or guidance condition. In this way, Bohm’s momentum does not need to be considered an external concept to the usual quantum theory; we find that this concept is indeed contained in a natural fashion within this theory [13]. Indeed, physically, it is a measure of the local spatial variations undergone by the phase, while the so-called Bohmian trajectories constitute, in turn, a local measure of the flux along precise streamlines, which have nothing to do with real trajectories followed by a quantum system [14, 15, 16].
@@
Following the above discussion, we can thus understand the so-called Bohmian trajectories either as hidden variables, providing us with an alternative perspective of quantum systems [17, 18, 19, 20, 6, 21, 22, 23], or, alternatively, as a methodological tool, worth applying to a wide variety of physical problems to get new insights [7, 8, 9, 24, 10], which may have either a quantum mechanical nature, or necessarily not. Here, we are going to focus on the second point of view, trying to settle in simple terms what a Bohmian trajectory is [15, 16] and what it can be good for [13], leaving aside expendable ontological connotations. This will also serve us to show how this quantum trajectory-based methodology can be extrapolated to other non-quantum contexts, which keep a certain resemblance at a formal level.
@@
To illustrate such claims, here two challenging wave phenomena in free space, directly associated with the phase, are analyzed and discussed by means of Bohmian mechanics: (1) solutions that contain an infinite amount of energy, with particular emphasis on the phenomenology of Airy wave packets due to their self-acceleration and shape-invariance properties [25], and (2) self-focusing solutions [26], i.e., solutions that converge to a singularity with time without the action of any external agent or nonlinear contribution [27, 28, 29, 30]. In both cases, it will be shown how, despite the peculiar nature of these solutions, they admit well-posed trajectory-based descriptions, which help us to understand in an intuitive manner the main features (the challenging ones) of their respective dynamical evolutions and, from them, the major role played in those dynamics by the local variations of the quantum phase. An interesting point here is thus the one concerning the symmetries displayed by the underlying velocity fields (or, equivalently, the associated phase fields), and hence the physical implications arising when such symmetries are broken.
@@
The work is organized as follows. In Sec. 2 a brief overview of the main theoretical aspects related to Bohmian mechanics is summarized, as well as its extension to standard paraxial wave optics, which relies on a Schrödinger-type equation, namely, the paraxial Helmholtz equation. Section 3 presents a discussion of the Airy beam propagation in terms of Bohmian trajectories, while spontaneous self-focusing is considered in Sec. 4. Finally, a series of concluding remarks are highlighted in Sec. 5.
@@ 2.1 Velocities and trajectories in standard quantum mechanics
Consider a spinless, nonzero mass quantum system described in the configuration space, at a time $t=0$, by the wave function $\psi({\bf r},0)$. If the system is in free space, the time evolution of the system wave function at any other later time is ruled by the simple time-dependent Schrödinger equation
@@
${\bf j}({\bf r},t)=\frac{\hbar}{2mi}\Big{[}\psi^{*}({\bf r},t)\nabla\psi({\bf r},t)-\psi({\bf r},t)\nabla\psi^{*}({\bf r},t)\Big{]}$ (3)
@@
and $\hat{\bf p}=-i\hbar\nabla$ is the usual momentum operator. Accordingly, the quantity ${\bf p}({\bf r},t)=m{\bf v}({\bf r},t)$ can be understood as a measure of the local value of the momentum, complementary of the corresponding expectation value, which provides us with a sort of average of ${\bf p}$ over the whole configuration space.
So far, all concepts introduced above are included within standard quantum mechanics and, therefore, there is no need to consider extra postulates. Let us now briefly overview the conventional Bohmian approach, which merges the standard quantum theory with the concept of hidden variable, i.e., assigning specific positions in configuration space to quantum systems [1, 2, 6, 31]. The evolution in time of these positions give rise to well-defined trajectories, known as Bohmian trajectories, which do not violate any fundamental law of quantum mechanics, such as the uncertainty relations or the complementarity principle. To reach such a trajectory-based formulation, Bohm considered [1, 6] the nonlinear transformation relation $\{\psi,\psi^{*}\}\to\{\rho,S\}$, from two complex field variables to two real field variables, such that
@@
and the same for $\psi^{*}$. After making the corresponding substitution into Eq. (1), and its conjugate, and then separating the real and imaginary parts of the resulting equation, one obtains the equations of motion for both $\rho({\bf r},t)$ and $S({\bf r},t)$, namely,
$\frac{\partial\rho({\bf r},t)}{\partial t}$ $=$ $-\nabla\cdot\left[\rho({\bf r},t)\frac{\nabla S({\bf r},t)}{m}\right],$ (6)
@@
Equation (6) corresponds to the continuity equation (2), although the velocity field appears recast in terms of the new phase variable, i.e., ${\bf v}=\nabla S/m$, while Eq. (7) is the so-called quantum Hamilton-Jacobi equation, which generalizes the classical version [6, 32] by including the extra contribution $Q({\bf r},t)$, known as Bohm’s potential or quantum potential,
@@
$\dot{\bf r}({\bf r},t)=\frac{{\bf p}({\bf r},t)}{m}=\frac{\nabla S({\bf r},t)}{m},$ (9)
where ${\bf p}$ is the so-called Bohm’s momentum, which corresponds to the quantity ${\bf p}=m{\bf v}$ introduced above. From a formal point of view, this equation of motion, known as the guidance condition or equation, tells us how a local velocity field, intrinsically associated with the spatial variations of the phase, rules the distribution of the probability density in configuration space. Maybe this is a rather challenging consequence in quantum mechanics. However, note that in optics the engineering of structured light beams [33], which precisely follow similar equations (see below), is actually possible by making light to pass through devices that produce such spatial phase variations (i.e., spatial light modulators). Even more, with the possibility to experimentally obtain weak measurements [34, 35], it has been possible to measure this transverse field and, from it, to infer the path followed by the corresponding trajectories [36, 37, 38] (which has nothing to do with the hypothetical pathways that could be expected to be followed by individual photons).
From a conventional point of view, we thus have a combination of standard tools in a hydrodynamic language (à la Madelung [5]) and a postulated momentum (à la Bohm [1]), which does not disprove the above pragmatic consideration of Bohmian mechanics as a valid methodology or quantum representation, beyond its commonly accepted conception of being a hidden-variable model. This leads us to a third aspect, which emphasizes the usefulness of this approach as a convenient computational tool. Specifically, when one deals with the issue of propagating the wave function following the Schrödinger equation, there are different strategies, the so-called spectral methods being a widely used methodology [39, 40, 41]. The basic idea of these methods is to recast the wave function as a linear superposition of eigenfunctions of the Laplacian, so that the nonlocal action of the kinetic term can be evaluated in a more accurate manner than with discretizing methods (where the accuracy of the method is associated with the number of neighbors included in the approximation of the discrete Laplacian). From a Bohmian point of view, this is a rather convenient strategy to obtain analytical expressions for the equation of motion in terms of the spectral or pseudo-spectral components of the wave function [9]. Indeed, if the method is based on the Fourier transform, these components can directly be the components of the discrete Fourier transform, which simplifies remarkably the numerical calculation of the trajectories and, at the same time, also increases their numerical accuracy (something very important in problems characterized by the presence and sudden emergence of nodes).
@@
where $c_{j}=|c_{j}|e^{\phi_{j}}$, ${\bf k}_{j}={\bf p}_{j}/\hbar$, and $E_{j}={\bf p}_{j}^{2}/2m$. Note that the ${\bf k}_{j}$ might correspond, for instance, to the discrete set of momenta that arise as soon as the configuration space is discretized in a mesh of equidistant knots, when we apply the fast Fourier method [42]. If Eq. (10) is substituted into Eqs. (4) or, equivalently, (9), we obtain the equation of motion
@@
2.2 From trajectories in standard quantum mechanics to rays in standard scalar optics
@@
Let us now assume that the energy propagates mostly along a given direction (say, $z$), while the effects of interest lie on perpendicular planes. This means that we can assume paraxial conditions, i.e., all important effects take place close to the $z$-axis as the field mainly moves forward parallel to it. Accordingly, we can write the time-independent field amplitude as $\Psi({\bf r}_{\perp},z)\approx\psi({\bf r}_{\perp},z)e^{ikz}$, i.e., as a plane wave accounting for the propagation along the $z$-axis, with a modulating amplitude along the transverse direction. With this assumption, and after neglecting slow second order derivatives, Eq. (15) becomes the paraxial Helmholtz equation,
@@
and $\nabla^{2}_{\perp}$ is replaced by $\nabla^{2}$, then we immediately recover Schrödinger’s equation for propagation in free space. Accordingly, considering the same transformation, from Eq. (9), we find that an analogous ray equation can also be obtained in optics, since
@@
The above relation basically says that the momentum and the wave vector are linearly related one another, since, making both sides to be equal, we obtain ${\bf p}/\hbar={\bf k}$, which is precisely de Broglie’s quantization condition, which dynamically associates the particle aspect (${\bf p}$) with a wave aspect (${\bf k}$ or, equivalently, $1/\lambda$) — also valid for photons if Einstein’s equations for the energy of a zero-mass particle and for the photon energy are made equal.
@@
where ${\bf v}_{\perp}$ is not a velocity in strict sense, but a field that behaves similarly in an “effective” manner. This effective field is analogous to the transverse momentum introduced in [45] to describe the properties of an atomic Mach-Zehnder interferometer. The integration of this quantity in $z$ provides us with trajectories that describe the spatial distribution of the electromagnetic energy or, equivalently, how the latter redistributes along the transverse direction as $z$ increases. It is worth noting that the outcomes experimentally reported in [37] correspond precisely to the transverse field (19), while the inferred trajectories are in compliance with what one would obtain by integrating this effective equation of motion. On the other hand, it is also worth emphasizing that Eq. (19) for paraxial conditions is an approximation to the general expression that arises from the Poynting vector [46].
@@
where ${\bf u}$ refers to the spatial coordinates (transverse, in the case of the paraxial optical system), $\xi$ is the propagation variable, and $\nabla^{2}_{\bf u}$ and $\nabla_{\bf u}$ are, respectively, the Laplacian and the gradient computed with respect to the spatial coordinates ${\bf u}$, and ${\bf v}_{\bf u}$ denotes the effective velocity field in the space defined by the coordinates ${\bf u}$. After integrating ${\bf v}_{\bf u}$ along the propagation variable $\xi$, for a set of initial conditions ${\bf u}_{0}(\xi=0)$, one obtains ensembles of trajectories parameterized in terms of the $\xi$, which provide us with an alternative description of the system propagation.